A normal distribution that has a mean of 100 and sd of 8. I have determined the z scores of x=18, 112, 98, 70, and 124. Now it is asking to determine the area with
p(x>108)
p(x<98)
p(92 p(84 p(x<84)
How to you calcute the answer to these. Please help

Answer by stanbon(75887)

A normal distribution that has a mean of 100 and sd of 8. I have determined the z scores of x=18, 112, 98, 70, and 124. Now it is asking to determine the area with
p(x>108)
z(108) = (108-100)/8 = 1
P(x > 108) = P(z > 1) = normalcdf(1,100) = 0.1587
-----
p(x<98)
z(98) = (98-100)/8 = -1/4
P(x < 98) = P(z < -1/4) = 0.4013
-------
p(x<84)
same procedure
====================
Cheers,
Stan H.
====================